OpenStudy (anonymous):

Dots sells a total of 265 T-shirts ($2) and shorts ($3). In April, total sales were $637. How many T-shirts and shorts did Dots sell?

OpenStudy (jdoe0001):

t = t-shirts sold s = shorts sold we know the t-shirts are $2 each, so if she sold "t" amount, the value of them will be 2*t or 2t we know the shorts are $3 each, so if she sold "s" amount, the value of them will be 3*s or 3s the total combined value amount is $637 thus one could say \(\bf 2t+3s=637\) we know she sold 265 total in quantity of "t" and "s" so one can say \(\bf t+s=265\) can you solve t+s=265 for "t"? what would that give?

OpenStudy (anonymous):

okay so all of that makes sense, but I am unsure of how to solve t+s=265

OpenStudy (jdoe0001):

well.. "isolate" t on the left-hand side what would that give you?

OpenStudy (anonymous):

265-s?

OpenStudy (jdoe0001):

yeap, so now we know that t = 265-s let us use that in our other equation, so \(\bf t+s=265\implies {\color{red}{ t}}=265-s \\ \quad \\ 2{\color{red}{ t}}+3s=637\implies 2({\color{red}{ 265-s}})+3s=637\) so there, solve for , or "isolate" "s" to find how many shorts she sold

OpenStudy (anonymous):

okay so i'm trying to figure it out: would i do 2x265 and 2xs which would look like: 530-2s+3s=637

OpenStudy (jdoe0001):

yeap thus \(\bf 530-2s+3s=637\implies 530+s=637\implies s=637-530\)

OpenStudy (anonymous):

okay great! so 107 shorts and 158 t-shirts, correct?

OpenStudy (jdoe0001):

okay great! so 107 shorts and 158 t-shirts, correct? \(\large \checkmark\)